OpenStudy (loser66):

show that if \[\frac{N_x-M_y}{xM-yN}=R\]where R depends on the quantity xy only. Then M+Ny'=0 has an integrating factor of the form\[\mu(xy)\]. Find a general formula for this integrating factor. Please, help.

OpenStudy (anonymous):

From definition for equation to have intgrating factor: \((M\mu)_y =( N\mu )_x\) differentiating \(M_y\mu+M\mu_y=N_x\mu+N\mu_x\) so \( \large \mu=\frac{M\mu_y-N\mu_x}{N_x-M_y}\) now take \(\huge \frac{1}{\mu}\) so if this quanitiy we call R and it depends only on xy, we see that 1/R=\(\mu\) which will also depend on xy

OpenStudy (loser66):

so, we can conclude that the form of that integrating factor is \[R= \frac {1}{\mu}\], right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just there is one point: R(xy) can be anything. For example: \(3(xy)^2+e^{xy}\) So \(R(xy)_{x}=R_xy\) and \(R(xy)_y=R_yx\) and you have to take this into the account for \(\mu\)

OpenStudy (anonymous):

you would check befor starting to look for integrating factor, :)

OpenStudy (loser66):

this problem is showing one. It means it is general case. I have another problem. They asked me apply this concept. I will carefully do it, If I get stuck, I tag you to get help. Thanks a lot.

OpenStudy (loser66):

My question is when Myko take \[R= \frac{1}{\mu}\] it doesn't form the form question asked for, I need x*M - y*N at denominator,

OpenStudy (loser66):

@miko @druminjosh . I have class now and have to go. Whenever you guys came online, please explain me steps. Thanks in advance.

OpenStudy (anonymous):

continuing with my 1º comment: \(\huge \frac{1}{\mu}=\frac{N_x-M_y}{M\mu_y-N\mu_x}\) this is already almost what your question states. Just need one more step. Now \(\mu(xy)\) is any function. So taking it's derivative respect to x and later also y will look something like this: \(\large\mu_y=\mu_{xy}x\) .This is just chain rule, if you wonder, :). Same for the other: \(\large\mu_x=\mu_{xy}y\). Now put this into the expretion for \(\huge \frac{1}{\mu}\) and rearange it a bit. \(\huge \frac{\mu_{xy}}{\mu}=\frac{N_x-M_y}{xM-yN}\) This is now exactly what you question states. So the relation between R and \(\mu\) is: \(\huge\frac{\mu_{xy}}{\mu}=R\) Hope it is ok now, :).

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

what @myko said. don't have anything to add, he beat me to it :) good job @myko (and thanks! save me a lot of work hehe)